OpenStudy (anonymous):

The gas piston of an M4 carbine rifle with in initial position of x0=5[mm].experiences a velocity defined by x’=3+5t for one second. What is the piston’s finale position?

OpenStudy (anonymous):

since x' is the velocity you need to integrate the expression 3+5t so as to get x, w/c is x = 5/2 x^2 +3t +5 where x is in mm and t in sec. Thus the final position is x(1) = 5/2 +3 +5 =21/2 mm

OpenStudy (anonymous):

sorry what is the meaning of x^2

OpenStudy (anonymous):

\[x^{2}\]

OpenStudy (anonymous):

Sorry my dear can you write it a little simple, i didn't get it, plz step by step im writing

OpenStudy (anonymous):

see, when we integrate 3+5t dt, it will be become 3t+5t square/2 + c, now how to find c

OpenStudy (anonymous):

Zekarais are you there

OpenStudy (anonymous):

\[x'=3+5t\]\[x =\int\limits(3+5t)dt\]\[x =\frac{ 5 }{ 2 }t^{2}+3t +c\]Since we told that the initial position is 5mm, we set t = 0 so as to get c, w/c is 5.

OpenStudy (anonymous):

oka after this

OpenStudy (anonymous):

Thus we will get this equation\[x =\frac{ 5 }{ 2 }t^{2}+3t+5\]

OpenStudy (anonymous):

after this equation, go ahead

OpenStudy (anonymous):

up to find the 21/0 ir 10.5 mm

OpenStudy (anonymous):

21/2 or 10.5

OpenStudy (anonymous):

To find the final position set t = 1, w/c is given.

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

how you found the C = 5

OpenStudy (anonymous):

what is w/c ?

OpenStudy (anonymous):

What was the position of the piston initially?

OpenStudy (anonymous):

5 mm,

OpenStudy (anonymous):

so it mean C= initial position ?

OpenStudy (anonymous):

5mm means the position at t = 0,right. yape

OpenStudy (anonymous):

after this step, x=52t2+3t+c how to find C, or we should put the 5 directly ?

OpenStudy (anonymous):

set t = 0, c= 5mm. It just a simple task.

OpenStudy (anonymous):

where you gone

OpenStudy (anonymous):

hmm.. please set t = 0 , c = 5 once

OpenStudy (anonymous):

Dear my friend the question said " if you put t = 0 you will get the position of the piston 5mm." I think this will help

OpenStudy (anonymous):

Sorry, I want to write complete step by step.

OpenStudy (anonymous):

I AM OUT.

OpenStudy (anonymous):

anyway thanks alot